The sum of Marian's age and David's age when they were married was 44. Five years after they were married, twice David's age minus three times Marian's age was -22. How old were each when they were married?

I have tried using the elimination method by using two equations, but I'm still not sure. Please help me with this.

Answer by mananth(16946)

when they married
Marian's age = x
David's age =y
x+y=44
After 5 years
Marian's age = x+5
David's age = y+5
2(y+5)-3(x+5)=-22
2y+10-3x-15=-22
2y-3x=-17
1 x + 1 y = 44 .............1
-3 x + 2 y = -14 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-2 x -2 y = -88
-3 x + 2 y = -14
Add the two equations
-5 x = -102.000
/ -5
x = 20 2/ 5 Marian's age
plug value of x in (1)
1 x + 1 y = 44
20.4 + 1 y = 44
1 y = 44 -20.4
1 y = 23.6
y = 23 3/ 5 David's age